OpenStudy (anonymous):

The length of a rectangle is 24 units. Can the perimeter x of the rectangle be 60 units when its width y is 11 units? No, the rectangle cannot have x = 60 and y = 11 because x = 48 + 2y No, the rectangle cannot have x = 60 and y = 11 because x = 24 + 2y Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 48 Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 24

OpenStudy (phi):

the perimeter of a rectangle is the sum of its sides. you have two sides, each 24 and 2 sides, each 11 24+24+11+11= 48+12 = 60 so the perimeter is 60

OpenStudy (phi):

but none of the answers make sense.